Two equal mass object experience a totally inelastic elastic collision. Mass 1 has an initial velocity of 10 m/s in the negative y-direction. Mass 2 has an initial velocity of 10 m/s in the positive x-direction. The collision occurs at the origin. What is the magnitude and direction of the velocity of the combined mass?
1
Expert's answer
2019-03-18T11:36:23-0400

"\overrightarrow{V_1} = 0*i -10*j"

"\overrightarrow{V_2} = 10*i +0*j"


"\overrightarrow{V_{final}} = \overrightarrow{V_1} + \overrightarrow{V_2} =0*i+10*i-10*j+0*j = 10*i - 10*j"


"|\overrightarrow{V_{final}}| = \sqrt{(10)^2 +(-10)^2} = 10\sqrt{2} \approx 14.1 m/s"


Answer: 14.1 m/s